OpenStudy (anonymous):

By what number must you multiply each side of the equation to produce the equivalent equation x = 10?

OpenStudy (anonymous):

EUATION ??????

OpenStudy (anonymous):

4/5 (1/5)x=8(1/5) 4/25x=8/5 4x=8/5(25) 4x=40 x=10 1/5 is correct